OpenStudy (anonymous):

Why can I not find the limit this way? See reply as, the question won't format my work correctly limit as x approaches positive infinity (x-x^(1/2)).

OpenStudy (anonymous):

whats your way?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty }x ^{1/2} +x\] \[\lim_{x \rightarrow \infty }x(1/x^{1/2}) + 1)\] \[\infty \times(0+1)\] 0

OpenStudy (anonymous):

limx→∞x-x^1/2 limx→∞x(1-1/x^(1/2)) ∞×(0+1) ∞

OpenStudy (anonymous):

nvm